Use the Laplace transform to solve the initial value problem: y - 3y + 2y = 4t + ezt, y(0) = 1, y(0) = -1
Answer #1

Salection : The given initial value problem is , - D y – 3y + 2y = 47+e31 yo)=1; y (6)= -1 | with A second order - linearUsing eqo , ③. and ③, and also LYtY = 1 and Ideat? = we have, SPY(5) - Syco) - yco)-34 3 Yis)- yCo)} + 2 y(s) = 4(1 + 1 S-3s4 - +53 +1352 +45 - 12_(Fachoning denaginator) s. 52 (5-3) CS-2) (S-1) factoring denominator we can write Y is) as, Tes) =collect the like terms s-753 +13.52 +45-12 3 S4 CATC+o+ E) +33(-6A +B-57-3D -46) +52 (11A - 62 +6C +20+36) +s(-6A +11B) -(68f Sawing all this A-3, B2 systems we have in Cari Dağ E = -2 .: from 299 2 | 54-733 + 1852 +45-12 = 3+ S2 (S-3) (S-2) (S-1) +